Use Euclidean algorithm to find integers u and v such that 149u + 139V = 1, where u and v are integers.
Answer #1

Solution :

Step one – use the Euclidean algorithm

(1) 149 = 1 × 139 + 10

(2) 139 = 13 × 10 + 9

(3) 10 = 1 × 9 + 1

(4) 9 = 9 × 1 + 0

(5)

A remainder of 0 has been obtained, so the gcd is the last non-zero remainder, hence

gcd(149,139) = 1 as required.

Step two – rearrange the earlier expressions to make the remainder the subject.

(3) 1 = 10 – 1 × 9

(2) 9 = 139 – 13 × 10

(1) 10 = 149 – 1 × 139

Step 3 – back substitute

Start with (3) 1 = 10 [× 1] – 1 × 9

Substitute in (2) 1 = 10 × 1 – 1(139 – 13 × 10)

1 = 10 × 1 – 139 × 1 + 13 × 10

1 = 10 × 14 – 139 × 1

Substitute in (1) 1 = (149 × 1 – 1 × 139) × 14 – 139 × 1

1 = 149 × 14 – 14 × 139 – 139 × 1

1 = 149 × 14 – 15 × 139

Thus,149u + 139v = 1 ;where u = 14 and v = –15.